# Minimizing readout-induced noise for early fault-tolerant quantum computers

### v2

@@ Abstract
Quantum error correcting code can diagnose potential errors and correct them based on measured outcomes by leveraging syndrome measurement. However, mid-circuit measurement has been technically challenging for early fault-tolerant quantum computers and the readout-induced noise acts as a main contributor to the logical infidelity. We present a different method for syndrome extraction, namely Generalized Syndrome Measurement, that requires only a single-shot measurement on a single ancilla, while the canonical syndrome measurement requires multiple measurements to extract the eigenvalue for each stabilizer generator. As such, we can detect the error in the logical state with minimized readout-induced noise. By adopting our method as a pre-check routine for quantum error correcting cycles, we can significantly reduce the readout overhead, the idling time, and the logical error rate during syndrome measurement. We numerically analyze the performance of our protocol using Iceberg code and Steane code under realistic noise parameters based on superconducting hardware and demonstrate the advantage of our protocol in the near-term scenario. As mid-circuit measurements are still error-prone for near-term quantum hardware, our method may boost the applications of early fault-tolerant quantum computing.
@@ I Introduction
Figure 1: Protocol scheme (a) Circuit scheme for canonical syndrome measurement (SM). A total number of $k$ (the number of stabilizer generators in the code) readouts are required to diagnose the error syndrome and apply proper correcting operations. While the ancilla qubit is being measured, all data qubits are subjected to decoherence due to idling. (b) Circuit scheme for (one-shot) generalized syndrome measurement (GSM). Only a single readout is needed to extract the partial information and detect the error, and the readout-induced idling error is minimized. (c) QEC cycle with generalized syndrome measurement. We divide the cycle into two stages: detection and correction. We first use the GSM method to check whether an error has occurred or not. If no error is detected, we skip the correction stage while ensuring that the logical state is in the codespace. If an error is detected, we follow the GSM routine by the canonical SM, extracting the full syndrome information and correcting the detected error. (d) A two-stabilizer code example illustrates the difference between the SM method and the GSM method. The value within the box denotes the measured eigenvalue for $S_{i}$ (SM) or the code-space projector $\bar{P}$ (GSM).
@@ II Protocol Description
Compared with the GSM method, the canonical SM method projects the input state to the codespace (+1 eigenspace of all $S_{i}$) by measuring the ancilla qubits sequentially. As each stabilizer $S_{i}$ requires an ancilla readout, at least $k$ readouts are required to tell whether the state is in the codespace or not and the readout-induced idling time will be $kt_{m}$ given $t_{m}$ is the time for a physical readout. The direct readout errors, like assignment errors and crosstalk errors, may also affect the fidelity of the logical state by $k$ times. In the GSM method, we require only a single measurement to extract partial syndrome information and diagnose if an error exists, at the cost of losing the specific information of which kind of error exists. Therefore, the GSM method can take the place of the SM method for quantum error detection (QED), where we post-select logical states based on readout outcomes. Fig. 1d provides an illustrative example of a code containing two stabilizer generators. For the purpose of QEC, the GSM method cannot completely replace the canonical SM method as all syndrome information is necessary. However, the GSM method can still bring benefits for QEC cycles: In a fault-tolerant quantum computing regime, The logical error rate of the input state is generally very low to ensure that we are under the noise threshold where logical qubits would outperform physical qubits. Therefore, the main purpose of a QEC cycle is to assure that no error has occurred in the input logical state, rather than revealing the specific information about the presented errors. Based on this assumption, we can integrate the GSM method in QEC cycles in an adaptive way and divide a cycle into two stages as presented in Fig. 1c: In the first stage, we detect if there is an error in the logical state using the GSM method. If no error is detected (in a more probable case), we assure the logical state is in the codespace, skip the next stage, and finish this QEC cycle. If at least an error is indeed detected (in a less probable case), we follow up with the canonical SM method to extract the full syndrome information and correct all detected errors. As long as the input state has sufficiently high fidelity, we can lower the average time of ancilla readouts by introducing the GSM method in such a combined way. For example, if the input state (encoded in a $k$-stabilizer code) is the ideal state with $99\%$ probability or a fully erroneous state with $1\%$ probability, the average readout-induced idling time for a noiseless QEC cycle with the GSM method can be approximated by $[0.99\cdot 1+0.01\cdot(1+k)]t_{m}=(1+0.01k)t_{m}$, which is much shorter than the time $kt_{m}$ for the canonical method even if $k=2$. As mid-circuit readouts are still very error-prone for current large-scale quantum hardware, we expect our method could significantly reduce the noise originating from readouts and benefit early fault-tolerant quantum applications.
@@
where $\{j\}_{i}$ is an index set associated with each subprojector and $\bigcup^{m}_{i=1}\{j\}_{i}=\{x\in\mathbb{Z}|0\leq x\leq k\}$. By measuring these subprojectors $\bar{P}_{i}$ separately, we can achieve an $m$-shot GSM protocol where the total readout time is still less than the canonical SM method. For example, a four-stabilizer code (e.g. the five-qubit $[[5,1,3]]$ code) with ${S_{1},S_{2},S_{3},S_{4}}$ can exploit a two-shot GSM by splitting the general projector into two subprojectors, one containing $S_{1}$ and $S_{2}$ while the other containing $S_{3}$ and $S_{4}$. We give a more detailed explanation in the supplemental material [28].
@@ III Gate implementation of $e^{i\pi\bar{P}}$
Figure 2: Gate Implementation of $e^{i\pi\bar{P}}$ (a) Sandwich implementation. $U_{\bar{P}}$ is the circuit decoding the code with projector $\bar{P}$ back to the trivial code state, and there is a multi-qubit control phase gate sandwiched in the middle between $U_{\bar{P}}$ and its conjugate $U^{*}_{\bar{P}}$. (b) The $e^{i\pi\bar{P}}$ can also be implemented by a sequence of control multi-qubit Pauli rotation gates.
Figure 3: Detecting the error with Iceberg code (a) Planar connectivity structure of Iceberg codes. All physical qubits are placed in an iceberg-like shape and the bottom/up qubit is connected to every other qubit. (b) The $[[4,2,2]]$ code with the required ancilla connectivity for GSM. (c) Decomposed circuit for GSM on the $[[4,2,2]]$ code. (d) Numerical simulation of four-qubit and six-qubit Iceberg code, with noise parameters of Sycamore [4] and ibm_brisbane processor [29]. The dashed curve indicates error rate when the gate control is perfect and all gate noise is gate-induced idling noise.
@@ IV Error detection with Iceberg code
which is the logical |+¯$\ket{\overline{+}}^{\otimes{2}}$ for $[[4,2,2]]$ code and |+¯$\ket{\overline{+}}^{\otimes{4}}$ for $[[6,4,2]]$ code up to local transversal Hadamard. Next, we apply a varying period of idling noise on every physical qubit to mimic realistic noise. We then separately use the canonical SM or the GSM method to detect the error and post-select the state without triggered syndrome. We adopt the sandwich implementation of the control projector gate and the decomposed circuit for the four-qubit code is shown in Fig. 3c, which requires the implementation of physical $CCZ$ gates or further decomposition. This implementation suits better for quantum hardware with less flexible qubit connectivity. The general implementation of the GSM method on the Iceberg code is discussed in our supplemental material [28]. We plot the logical error rate against the idling time in Fig. 3d as well as the raw physical state prepared in |$\ket{+}^{\otimes 2m}$ under idling noise for reference. We also plot the logical error rate when the gate noise is fully dominated by gate-induced decoherence as the dashed curve with the same colors.
@@ V Error correction with Steane code
Figure 4: Correcting the error using seven-qubit Steane code (a) Planar connectivity structure of Steane code for the GSM method. On each color plaquette, the ancilla qubit is only required to connect two neighboring physical qubits on the edge. (b) Circuit scheme for three-shot GSM on Steane code. The check on each plaquette is performed sequentially but the ancilla qubits are measured in the same time. (c) Detailed implementation for plaquette check. (d) Logical error rate under idling noise after performing QEC cycle with (without) GSM and the corresponding number of average readout required. The inset figure is the logical error rate after just performing post-selection with quantum error detection (y axis label in the unit of 0.001).
@@
We use a projective noise parameter set for Steane code simulation because the state-of-the-art quantum hardware is still too noisy for logical qubits to outperform physical qubits [28]. We initialize the logical qubit in $\ket{\bar{+}}$ with noisy circuits, apply idling noise to every physical qubit, and perform quantum error correction on the noisy logical state. We compare the performance of the QEC cycle with GSM (Fig. 1c) and without GSM (Fig. 1a), and plot the logical error rate and the average number of readout in Fig. 4d. More simulation details can be found in the Method section as well.
@@ VI Discussion
A natural question arise from our work is: Does there exist a range where the GSM method doesn’t bring advantage on the logical fidelity? As the GSM method trades additional gate overhead for reduced readout overhead, the ratio between readout time and gate time can be used as the criterion. We numerically study the effect of readout time and gate time on the relative performance of the SM method and the GSM method in the supplemental material [28] and find out the parameter of practical superconducting quantum hardware falls in the regime where the GSM method is advantageous. Therefore, unless there is a technical breakthrough that shortened the physical readout time drastically, we believe the GSM method will bring practical advantage for near-term quantum applications.
Besides using the GSM method as a pre-check routine, fault-tolerant quantum computing can also benefit from GSM by exploit concatenated codes and treat detected physical errors as logical erasure errors [38, 28], which doesn’t require explicit quantum error correction in the inner code but only need to correct logical erasure in the outer loss-tolerant codes. In such a case, our GSM method can be exploited to improve the robustness against readout-induced error.
Notably, as the $e^{i\pi\bar{P}}$ gate used in the GSM method needs to be decomposed into elementary gates, local errors might propagate to other physical qubits and produce correlated errors. However, as long as readout-induced idling error, which is a global noise channel, still serves as the main noise contributor in the near-term scenario, the correlated errors won’t weigh over the readout-induced errors and pose a critical threat to the validness of our proposed method. Moreover, we may also use the idea of Shor-style syndrome measurement, which rely on pre-entangled ancilla qubits and native multi-qubit Pauli gates to achieve gate-level fault tolerance at the cost of more ancilla qubits and time overhead [28].
@@
Acknowledgement

### v3

@@ Abstract
Quantum error correcting code can diagnose potential errors and correct them based on measured outcomes by leveraging syndrome measurement. However, mid-circuit measurement has been technically challenging for early fault-tolerant quantum computers and the readout-induced noise acts as a main contributor to the logical infidelity. We present a different method for syndrome extraction, namely Generalized Syndrome Measurement, that requires only a single-shot measurement on a single ancilla, while the canonical syndrome measurement requires multiple measurements to extract the eigenvalue for each stabilizer generator. As such, we can detect the error in the logical state with minimized readout-induced noise. By adopting our method as a pre-check routine for quantum error correcting cycles, we can significantly reduce the readout overhead, the idling time, and the logical error rate during syndrome measurement. We numerically analyze the performance of our protocol using Iceberg code and Steane code under realistic noise parameters based on superconducting hardware and demonstrate the advantage of our protocol in the near-term scenario. As mid-circuit measurements are still error-prone for near-term quantum hardware, our method could boost the applications of early fault-tolerant quantum computing.
@@ I Introduction
Figure 1: (a) Circuit scheme for canonical syndrome measurement (SM). A total number of $k$ (the number of stabilizer generators in the code) readouts are required to diagnose the error syndrome and apply proper correcting operations. While the ancilla qubit is being measured, all data qubits are subjected to decoherence due to idling. (b) Circuit scheme for (one-shot) generalized syndrome measurement (GSM). Only a single readout is needed to extract the partial information and detect the error, and the readout-induced idling error is minimized. (c) QEC cycle with generalized syndrome measurement. We divide the cycle into two stages: detection and correction. We first use the GSM method to check whether an error has occurred or not. If no error is detected, we skip the correction stage while ensuring that the logical state is in the codespace. If an error is detected, we follow the GSM routine by the canonical SM, extracting the full syndrome information and correcting the detected error. (d) A two-stabilizer code example illustrates the difference between the SM method and the GSM method. The value within the box denotes the measured eigenvalue for $S_{i}$ (SM) or the code-space projector $\bar{P}$ (GSM).
@@ II Protocol Description
Compared with the GSM method, the canonical SM method projects the input state to the codespace (+1 eigenspace of all $S_{i}$) by measuring the ancilla qubits sequentially. As each stabilizer $S_{i}$ requires an ancilla readout, at least $k$ readouts are required to tell whether the state is in the codespace or not and the readout-induced idling time will be $kt_{m}$ given $t_{m}$ is the time for a physical readout. The direct readout errors, like assignment errors and crosstalk errors, may also affect the fidelity of the logical state by $k$ times. In the GSM method, we require only a single measurement to extract partial syndrome information and diagnose if an error exists, at the cost of losing the specific information of which kind of error exists. Therefore, the GSM method can take the place of the SM method for quantum error detection (QED), where we post-select logical states based on readout outcomes. Fig. 1d provides an illustrative example of a code containing two stabilizer generators. For the purpose of QEC, the GSM method cannot completely replace the canonical SM method as all syndrome information is necessary. However, the GSM method can still bring benefits for QEC cycles: In a fault-tolerant quantum computing regime, The logical error rate of the input state is generally very low to ensure that we are under the noise threshold where logical qubits would outperform physical qubits. Therefore, the main purpose of a QEC cycle is to assure that no error has occurred in the input logical state, rather than revealing the specific information about the presented errors. Based on this assumption, we can integrate the GSM method in QEC cycles in an adaptive way and divide a cycle into two stages as presented in Fig. 1c: In the first stage, we detect if there is an error in the logical state using the GSM method. If no error is detected (in a more probable case), we assure the logical state is in the codespace, skip the next stage, and finish this QEC cycle. If at least an error is indeed detected (in a less probable case), we follow up with the canonical SM method to extract the full syndrome information and correct all detected errors. As long as the input state has sufficiently high fidelity, we can lower the average time of ancilla readouts by introducing the GSM method in such a combined way. For example, if the input state (encoded in a $k$-stabilizer code) is the ideal state with $99\%$ probability or a fully erroneous state with $1\%$ probability, the average readout-induced idling time for a noiseless QEC cycle with the GSM method can be approximated by $[0.99\cdot 1+0.01\cdot(1+k)]t_{m}=(1+0.01k)t_{m}$, which is much shorter than the time $kt_{m}$ for the canonical method even if $k=2$.
Notably, the idea of adaptive syndrome measurement has already been used for Shor-style error correction [28, 29], but we are using it here in a different way: The first measurement is for the whole codespace projector, and the following measurements are for each stabilizer generator. Our method could also be further made in Shor-style fault-tolerant way (see Appendix F). As mid-circuit readouts are still very error-prone for current large-scale quantum hardware, we expect our method could significantly reduce the noise originating from readouts and benefit early fault-tolerant quantum applications.
@@
where $\{j\}_{i}$ is an index set associated with each subprojector and $\bigcup^{m}_{i=1}\{j\}_{i}=\{x\in\mathbb{Z}|0\leq x\leq k\}$. By measuring these subprojectors $\bar{P}_{i}$ separately, we can achieve an $m$-shot GSM protocol where the total readout time is still less than the canonical SM method. For example, a four-stabilizer code (e.g. the five-qubit $[[5,1,3]]$ code) with ${S_{1},S_{2},S_{3},S_{4}}$ can exploit a two-shot GSM by splitting the general projector into two subprojectors, one containing $S_{1}$ and $S_{2}$ while the other containing $S_{3}$ and $S_{4}$. We give a more detailed explanation in the Appendix B.
@@ III Gate implementation of $e^{i\pi\bar{P}}$
Figure 2: Gate Implementation of $e^{i\pi\bar{P}}$. (a) Sandwich implementation. $U_{\bar{P}}$ is the circuit decoding the code with projector $\bar{P}$ back to the trivial code state, and there is a multi-qubit control phase gate sandwiched in the middle between $U_{\bar{P}}$ and its conjugate $U^{*}_{\bar{P}}$. (b) The $e^{i\pi\bar{P}}$ can also be implemented by a sequence of control multi-qubit Pauli rotation gates.
Figure 3: (a) Planar connectivity structure of Iceberg codes. All physical qubits are placed in an iceberg-like shape and the bottom/up qubit is connected to every other qubit. (b) The $[[4,2,2]]$ code with the required ancilla connectivity for GSM. (c) Decomposed circuit for GSM on the $[[4,2,2]]$ code. (d) Numerical simulation of four-qubit and six-qubit Iceberg code, with noise parameters of Sycamore [4] and ibm_brisbane processor [30]. The dashed curve indicates error rate when the gate control is perfect and all gate noise is gate-induced idling noise.
@@ IV Error detection with Iceberg code
which is the logical |+¯$\ket{\overline{+}}^{\otimes{2}}$ for $[[4,2,2]]$ code and |+¯$\ket{\overline{+}}^{\otimes{4}}$ for $[[6,4,2]]$ code up to local transversal Hadamard. Next, we apply a varying period of idling noise on every physical qubit to mimic realistic noise. We then separately use the canonical SM or the GSM method to detect the error and post-select the state without triggered syndrome. We adopt the sandwich implementation of the control projector gate and the decomposed circuit for the four-qubit code is shown in Fig. 3c, which requires the implementation of physical $CCZ$ gates or further decomposition. This implementation suits better for quantum hardware with less flexible qubit connectivity. The general implementation of the GSM method on the Iceberg code is discussed in Appendix E. We plot the logical error rate against the idling time in Fig. 3d as well as the raw physical state prepared in |$\ket{+}^{\otimes 2m}$ under idling noise for reference. We also plot the logical error rate when the gate noise is fully dominated by gate-induced decoherence as the dashed curve with the same colors.
@@ V Error correction with Steane code
Figure 4: (a) Planar connectivity structure of Steane code for the GSM method. On each color plaquette, the ancilla qubit is only required to connect two neighboring physical qubits on the edge. (b) Circuit scheme for three-shot GSM on Steane code. The check on each plaquette is performed sequentially but the ancilla qubits are measured in the same time. (c) Detailed implementation for plaquette check. (d) Logical error rate under idling noise after performing QEC cycle with (without) GSM and the corresponding number of average readout required. The inset figure is the logical error rate after just performing post-selection with quantum error detection (y axis label in the unit of 0.001).
@@
We use a projective noise parameter set for Steane code simulation because the state-of-the-art quantum hardware is still too noisy for logical qubits to outperform physical qubits see Appendix C. We initialize the logical qubit in $\ket{\bar{+}}$ with noisy circuits, apply idling noise to every physical qubit, and perform quantum error correction on the noisy logical state. We compare the performance of the QEC cycle with GSM (Fig. 1c) and without GSM (Fig. 1a), and plot the logical error rate and the average number of readout in Fig. 4d. More simulation details can be found in the Method section as well.
@@ VI Discussion
A natural question arise from our work is: Does there exist a range where the GSM method doesn’t bring advantage on the logical fidelity? As the GSM method trades additional gate overhead for reduced readout overhead, the ratio between readout time and gate time can be used as the criterion. We numerically study the effect of readout time and gate time on the relative performance of the SM method and the GSM method in the supplemental material in Appendix D and find out the parameter of practical superconducting quantum hardware falls in the regime where the GSM method is advantageous. Therefore, unless there is a technical breakthrough that shortened the physical readout time drastically, we believe the GSM method will bring practical advantage for near-term quantum applications.
Besides using the GSM method as a pre-check routine, fault-tolerant quantum computing can also benefit from GSM by exploit concatenated codes and treat detected physical errors as logical erasure errors [39], which doesn’t require explicit quantum error correction in the inner code but only need to correct logical erasure in the outer loss-tolerant codes (see Appendix G). In such a case, our GSM method can be exploited to improve the robustness against readout-induced error.
Notably, as the $e^{i\pi\bar{P}}$ gate used in the GSM method needs to be decomposed into elementary gates, local errors might propagate to other physical qubits and produce correlated errors. However, as long as readout-induced idling error, which is a global noise channel, still serves as the main noise contributor in the near-term scenario, the correlated errors won’t weigh over the readout-induced errors and pose a critical threat to the validness of our proposed method. Moreover, we may also use the idea of Shor-style syndrome measurement, which rely on pre-entangled ancilla qubits and native multi-qubit Pauli gates to achieve gate-level fault tolerance at the cost of more ancilla qubits and time overhead. (see Appendix F)
@@ Appendix A Simulation details
Appendix B $m$-shot Generalized Syndrome Measurement with subprojectors
Figure 5: Circuit scheme for $m$-shot generalized syndrome measurement.
In the main text, we introduced the generalized syndrome measurement (GSM) which only requires a single readout to detect the error. The GSM method requires the ability to implement controlled projector gates $e^{i\pi\bar{P}}$, which could bring large computational overhead when decomposed into elementary gates and the connectivity requirement could be very strict. Here we introduce the modified method by exploiting a bit more ancilla qubits, which could help reach a smooth trade-off between the gate complexity/connectivity and the readout complexity in the GSM method.
For code $Q$ with stabilizer generator $\{S_{i}\}$, the general projector for one-shot GSM method is
$\bar{P}=\prod^{k}_{i}P_{i}=\frac{1}{2^{k}}\prod^{k}_{i}(I+S_{i})=\frac{1}{K}\sum^{K}_{i}M_{i}.$ (11)
As we showed in the main text, we can use a control-$e^{i\pi\bar{P}}$ gate and a single-shot measurement on the ancilla qubit to act the projector on the noisy state nondeterministically. Instead of using only one projector, we can group several subprojectors $\bar{P}_{i}$ such that their product is still $\bar{P}$:
$\bar{P}=\prod^{m}_{i}\bar{P}_{i},$ (12)
where
$\bar{P}_{i}=\prod_{\{j\}_{i}}P_{j}$ (13)
and $\{j\}_{i}$ is an index set associated with each subprojector and $\bigcup^{m}_{i=1}\{j\}_{i}=\{x\in\mathbb{Z}|0\leq x\leq k\}$. As we have $m$ subprojectors, we need $m$ ($1\leq m\leq k$) readouts for partial syndrome extraction here. We can use the circuit show in Fig. 5 to extract the partial syndrome in a QEC cycle.
As a convenience choice, we can choose to group stabilizers pairwise as $\{S_{2i-1},S_{2i}\}$, which allows us to define the two-element subprojectors $\bar{P}^{\textbf{2}}_{i}$
$\bar{P}^{\textbf{2}}_{i}=P_{2i-1}P_{2i}=\frac{1}{4}(I+S_{2i-1})(I+S_{2i}),$ (14)
and it’s easy to see
$\bar{P}=\prod^{k/2}_{i}\bar{P}^{\textbf{2}}_{i}.$ (15)
In such a way, we can halve the amount of readouts while the implementation of the projector gates $e^{i\pi\bar{P}^{\textbf{2}}_{i}}$ is still not too complicated. For instance, the [[5, 1, 3]] code has four stabilizer generators
$\begin{split}&S_{1}=ZXXZI\quad S_{2}=IZXXZ\\ &S_{3}=ZIZXX\quad S_{4}=XZIZX,\end{split}$ (16)
and we can just use a single readout to measure the projector
$\bar{P}=\frac{1}{16}(I+ZXXZI)(I+IZXXZ)\\ (I+ZIZXX)(I+XZIZX).$ (17)
This would require 18 two-qubit gates (depth 12) and a five-qubit Toffoli gate in total to implement the projector gate. Rather if we measure the following two subprojectors
$\begin{split}&\bar{P}_{1}=\frac{1}{4}(I+ZXXZI)(I+IZXXZ)\qquad\\ &\bar{P}_{2}=\frac{1}{4}(I+ZIZXX)(I+XZIZX),\end{split}$ (18)
, we only need 12 two-qubit gates (depth 6) and a Toffoli gate for each subprojector gate and the overall gate overhead is 24 two-qubit gates (depth 12) and two Toffoli gates.
Figure 6: Adoption of the GSM method on scalable two-dimensional color codes (a) a QEC cycle with the GSM method. Yellow dashed circles represent for the local GSM routine on each plaquette. When error rate is low, we will mostly end with just a single-shot readout time. We follow with local stabilizer check when the location of error(s) is detected to correct the error(s), and the time consumed is two-shot readout time. (b) a canonical QEC cycle without the GSM method. Whether an error is detected or not, the time required is always two-shot readout time.
Importantly, the partial syndrome information extracted with the GSM method can be further used in the correction stage in the QEC cycle. We only need to perform canonical syndrome measurement for stabilizers associated with triggered subprojectors, and stabilizers associated with untriggered subprojectors doesn’t need to be measured again. Taking the two-shot GSM method on the $[[5,1,3]]$ code as an example: If only the second subprojector $\bar{P}_{2}$ is triggered, we only need to perform canonical syndrome measurement for stabilizer $ZIZXX$ and $XZIZX$, and there is no need to measure $ZXXZI$ and $IZXXZ$ again as we already assured that the logical state is in the space of subprojector $\bar{P}_{1}$.
For another example, the two-dimensional color code [43] has stabilizer $\{X^{\otimes n},Z^{\otimes n}\}$ supported on each $n$-edge plaquette. We can use a subprojector associated with each plaquette
$\bar{P}_{plaq}=\frac{1}{4}(I+X^{\otimes n})(I+Z^{\otimes n})$ (19)
to extract partial syndrome information (Fig.  6). For this kind of code, a QEC cycle with our GSM method only takes a single readout cycle when no error occurs, and at most two readout cycles in total when at least an error occurs. To be compared with, a canonical QEC cycle without the GSM method always takes two readout cycles. Therefore, our GSM method can always reduce the time of readout for two-dimensional color code, providing a scalable option for fault-tolerant quantum computing.
In general, there is no restriction on how to group the stabilizer generators into subprojectors, and the subprojectors can be chosen to mostly suit the connectivity and practical limitation of quantum hardware. The number of readouts can be chosen smoothly from 1, corresponding to the one-shot GSM method proposed in the main text with the largest gate overhead (minimum readout overhead), to $k$, corresponding to the canonical syndrome measurement with the minimum gate overhead (largest readout overhead).
Appendix C Steane code simulation with state-of-the-art noise parameter
Figure 7: Steane code simulation with state-of-the-art noise parameter.
In the main text, we present numerical results for the GSM method on the Steane code with projective noise parameters. The main reason why we don’t stick with state-of-the-art noise parameter is the current superconducting quantum hardware is still too noisy and not in the regime where logical qubit would outperform physical qubit, as shown in Fig. 7. The logical error rate after the encoding and decoding is much larger than the physical error rate, which makes QEC not really meaningful in such a regime. However, we still see some advantage of the GSM method over the canonical SM method when the input logical state doesn’t suffer too long idling time. As longer idling noise is applied to the logical state, the average number of readout of a QEC cycle with the GSM method increases over two (canonical SM method) and the GSM method start to lose its advantage. As raw physical qubits dominates over the logical qubits, this regime doesn’t post a critique to our proposed GSM method.
Appendix D benchmark the advantageous regime of the GSM method
To understand the advantageous regime of the GSM method, we ramp both the readout time and the gate time for both the Sycamore and the ibm_brisbane setting within a practical range that might be reached in the future, while keep other noise parameters fixed. We compare the performance for both the GSM and the canonical SM method and plotted the fidelity difference $F_{gsm}-F_{sm}$ regarding the gate time and the readout time. We simulated both the $[[4,2,2]]$ code in Fig. 8 and the $[[6,4,2]]$ code in Fig. 9 to explore the condition when the GSM method is advantageous. The advantageous regime for the GSM(SM) method is plotted with red(blue) area. From the simulated results, we see the state-of-the-art parameter is in a range where the GSM method will beat the canonical SM method. Unless the readout duration is significantly reduced in the scalable quantum hardware, we will still stay in the advantageous regime for the GSM method.
Figure 8: Fidelity difference $F_{gsm}-F_{sm}$ with $[[4,2,2]]$ code depending on various readout time and gate time when other noise parameters are fixed. The yellow stars denote where the state-of-the-art operation time is.
Figure 9: Fidelity difference $F_{gsm}-F_{sm}$ with $[[6,4,2]]$ code depending on various readout time and gate time when other noise parameters are fixed. The yellow stars denote where the state-of-the-art operation time is.
Appendix E Projector gate implementation for Iceberg code
Although the implementation of the projector gate is not unique, here we present an implementation for general Iceberg code in Fig. 10. In such an implementation, the 2Q gate depth is only $4m$. Comparatively, the overall 2Q depth of the canonical syndrome measurement is $44m+$ 4.
Figure 10: the projector gate implementation of for $[[2m+2,2m,2]]$ Iceberg code.
Appendix F Fault-tolerance in analog with Shor-style syndrome measurement
In the long term, we have to consider the gate-level fault tolerance of our method to avoid the impact of correlated errors. To achieve this in our method, we require the quantum hardware to be able to implement native multi-qubit (control) Pauli phase gates and does not rely on decomposition into elementary gates. It’s reasonable to make such an assumption as various quantum hardware has demonstrated potential for native multi-qubit gates [12, 42]. We also demand that the native global gates don’t suffer from inner correlated errors. As such, we only need to consider single Pauli errors induced by the gate operations up to the main order.
By using the idea of Shor-style syndrome measurement, i.e. preparing entangled ancilla states, we can achieve a fault-tolerant circuit for our method. We plotted the GSM circuit for [4, 2, 2] code as an example in Fig. 11. The entangled ancilla states can prevent errors occurred in any single ancilla qubit. For errors occurred after the multi-qubit Pauli gates, they will either be detected by the following Pauli gates or commute with the following Pauli gates and will not be propagated to other qubits (Fig. 11).
Importantly, we note this fault tolerance cannot be achieved if we decompose the gate into elementary gates (two-qubit gates). That’s because each Pauli phase gate is mediating global interaction and an error occurred after a decomposed gate could propagate to every qubit at the end of the Pauli phase gate. In early fault-tolerant setting, this may not be a severe problem if the gain from the reduced readout-induced noise is good enough, as shown in our simulation results.
Figure 11: Generalized Syndrome Measurement in analog with Shor-style syndrome measurement for [[4, 2, 2]] code. Ancilla qubits (green) are prepared in the GHZ state (|$(\ket{000}+\ket{111})/\sqrt{2}$ and each ancilla qubit controls a multi-qubit Pauli gate. An error occurred after the gates will either be detected (left) or commute with the following gates and will not propagate (right).
Appendix G Quantum Error Correction with logical erasure conversion
We can exploit the benefit of concatenated quantum codes to correct the errors indirectly (Fig. 12). The idea is initially proposed in [39] for cluster state generations. The inner codes are set to be small-size quantum code used to detect the physical errors, and the outer codes are used to correct the loss (erasure) errors. Whenever we detect an error in the inner level, we actively treat it as a logical erasure error in the outer level. As outer codes are loss-tolerant, we can restore the logical information in the presence of logical erasures and thus correct the physical errors indirectly. We do not need the syndrome information and correct the Pauli errors directly, and the GSM method can help reduce the readout-induced infidelity in such a scenario.
Figure 12: Scheme for concatenated codes with logical erasure conversion. In such a scenario, the ability to detect the errors is sufficient to correct physical errors.